# Learnability Window in Gated Recurrent Neural Networks

### 20 Mar 2026 v7

@@ Appendix C Envelope validation
As discussed in Section 5.1, the envelope $f(\ell)$ defined in (8) relies on transport factors $\Gamma^{(q)}_{t,\ell}$ obtained from the first-order diagonal expansion of the Jacobian product $\mathcal{M}_{t,\ell}$. This expansion is what makes it possible to assign each hidden unit a lag-dependent rate $\mu^{(q)}_{t,\ell}$ with a transparent decomposition into gating and mixing contributions—a structure that cannot be read off the diagonal of $\mathcal{M}_{t,\ell}$ directly. This appendix verifies that the interpretability gained by the first-order expansion comes at no cost to the envelope’s decay profile: $f(\ell)$ preserves the decay trend of the exact-transport envelope $f_{\mathcal{M}}(\ell)=\sum_{q=1}^{H}|\Lambda^{(q)}[\mathcal{M}_{t,\ell}]_{qq}|$. We establish this with model-free linear statistics that assume no parametric form for the decay.
@@
[…]
@@ Results.
Spearman $\rho\geq 0.979$ across all 75 runs confirms that the first-order expansion never disrupts the lag ordering of the envelope. Pearson $r\geq 0.82$ (mean $0.96$) on the $\log_{10}$-transformed envelopes establishes that $\log f\approx a\,\log f_{\mathcal{M}}+b$, i.e. the two envelopes are related by a power-law rescaling $f\propto f_{\mathcal{M}}^{\,a}$. Such a rescaling preserves the monotonic and scaling structure on which the learnability analysis depends.
@@ Appendix D Projected adaptive base rate via Rayleigh quotient
Appendix E Noise floor of the empirical matched statistic
@@
This appendix clarifies the assumptions underlying the noise model used in the manuscript and justifies the scaling of stochastic fluctuations in the empirical matched statistic. The key distinction is between exact finite-sample distributions and domain-of-attraction arguments, which suffice to characterize scaling behavior. Throughout, we rely on the linearity of the matched statistic and on standard properties of heavy-tailed random variables [42].
By construction, the empirical matched statistic at lag $\ell$ is an average over $N$ independent training sequences,
$\widehat{S}_{N}(\ell)=\frac{1}{N}\sum_{n=1}^{N}S^{(n)}_{t,\ell}$,
where $S^{(n)}_{t,\ell}$ denotes the statistic computed from the $n$th sequence. Independence across sequences implies that the summands are independent for fixed $(t,\ell)$. Empirical studies of stochastic gradients in deep neural networks consistently report heavy-tailed fluctuations, typically with tail index $\alpha\in(1,2)$. Accordingly, we assume that the centered per-sequence statistic $S^{(n)}_{t,\ell}-\mathbb{E}[S_{t,\ell}]$ belongs to the domain of attraction of a symmetric $\alpha$-stable law. This assumption does not require exact $\mathcal{S}\alpha\mathcal{S}$ behavior at finite sample sizes; it only requires regularly varying tails with index $\alpha$.
The matched statistic $S_{t,\ell}$ is a finite linear combination of gradient components (Eq. (13)). Linearity preserves the tail index of heavy-tailed variables, so $S_{t,\ell}-\mathbb{E}[S_{t,\ell}]$ also lies in the domain of attraction of a symmetric $\alpha$-stable law, with a lag-dependent scale parameter $\sigma_{\alpha}(\ell)$. For analytical convenience, we represent this domain of attraction by a canonical $\mathcal{S}\alpha\mathcal{S}(0,\sigma_{\alpha}(\ell))$ random variable. Stability under summation then implies
@@
$\widehat{S}_{N}(\ell)-\mathbb{E}[S_{t,\ell}]\sim\mathcal{S}\alpha\mathcal{S}\!\left(0,\sigma_{\alpha}(\ell)N^{1/\alpha-1}\right)=\mathcal{S}\alpha\mathcal{S}\!\left(0,\frac{\sigma_{\alpha}(\ell)}{N^{\,1-1/\alpha}}\right).$
@@
Note that the $N^{-1/\alpha}$ scaling applies to sums of $\alpha$-stable variables, whereas the empirical average introduces an additional factor $1/N$, yielding the rate $N^{-(1-1/\alpha)}$. Thus, the typical magnitude of stochastic fluctuations decays as $N^{-(1-1/\alpha)}$. For $\alpha=2$ this recovers the classical Gaussian rate $N^{-1/2}$, while for $\alpha<2$ concentration is slower due to heavy-tailed noise. This noise floor determines the detectability thresholds and sample-complexity bounds derived in the manuscript and underlies the compression of the learnability window in the heavy-tailed regime.
@@ Appendix F LAN-based KL bound for averaged $\alpha$-stable location models
For completeness, we derive the KL lower bound used in Section 5.4.2 for the empirical matched statistic $\widehat{S}_{N}(\ell)$, which is defined as an average over $N$ independent training sequences. Unlike the classical LAN setting for product measures, the present analysis concerns a shrinking-scale $\alpha$-stable location family. The argument follows the general LAN framework for triangular arrays of statistical experiments with vanishing noise levels [52].
@@
We consider the statistical experiment induced by the empirical matched statistic $\widehat{S}_{N}(\ell)$ at a fixed lag $\ell$, which is defined as an average over $N$ independent training sequences. As shown in Appendix E, under heavy-tailed gradient fluctuations with tail index $\alpha>1$, the centered statistic $\widehat{S}_{N}(\ell)-\mathbb{E}[S_{t,\ell}]$ exhibits stochastic fluctuations whose typical magnitude scales as $N^{-(1-1/\alpha)}$.
To formalize hypothesis testing at lag $\ell$, we model the distribution of the centered empirical statistic by a location family with shrinking scale. Specifically, let $Z$ be a symmetric $\alpha$-stable random variable with unit scale and characteristic exponent $\alpha>1$, and define the observation
@@
Here, $Y_{N}$ represents the random fluctuation of the empirical matched statistic, and $\theta\in\mathbb{R}$ is a location parameter encoding the presence or absence of a lag-$\ell$ dependency.
@@
For $\alpha>1$, symmetric $\alpha$-stable laws are absolutely continuous and admit a density. Denoting by $P_{\theta,N}$ the law of $Y_{N}$, its density is given by
@@
where $f_{0}$ is the density of a fixed symmetric $\alpha$-stable distribution with unit scale.
@@
For $\alpha>1$, symmetric $\alpha$-stable densities possess finite Fisher information for the location parameter [23, 34, 42, 52]. As $s_{N}\to 0$, the family $\{P_{\theta,N}\}_{N\geq 1}$ forms a triangular array of location models that is locally asymptotically normal with normalization rate
@@
In the main text, the detection and non-detection hypotheses correspond to two symmetric location parameters separated by a fixed amount $\Delta(\ell)$:
$\theta_{1}=+\tfrac{1}{2}\Delta(\ell),\qquad\theta_{0}=-\tfrac{1}{2}\Delta(\ell),$
so that $\theta_{1}-\theta_{0}=\Delta(\ell)$.
The empirical matched statistic has a shrinking noise scale $s_{N}=\sigma_{\alpha}(\ell)N^{1/\alpha-1}$ (Appendix E). To express the fixed separation in the LAN framework, we measure it in units of this scale and define the normalized separation
@@ Appendix G From KL divergence to mutual information and Fano bounds
Fix a lag $\ell$. The hypotheses in Sec. 5.4.2 correspond to two $\alpha$–stable location models for the empirical matched statistic $S=\widehat{S}_{N}(\ell)$, obtained by averaging over $N$ independent training sequences. Introduce a binary label

### 28 Mar 2026 v8

@@ Appendix C Envelope validation
As discussed in Section 5.1, the envelope $f(\ell)$ defined in (8) relies on transport factors $\Gamma^{(q)}_{t,\ell}$ obtained from the first-order diagonal expansion of the Jacobian product $\mathcal{M}_{t,\ell}$. This expansion is what makes it possible to assign each hidden unit a lag-dependent rate $\mu^{(q)}_{t,\ell}$ with a transparent decomposition into gating and mixing contributions—a structure that cannot be read off the diagonal of $\mathcal{M}_{t,\ell}$ directly. This appendix verifies that the interpretability gained by the first-order expansion comes at no cost to the envelope’s decay profile: $f(\ell)$ preserves the decay trend of $f_{\mathcal{M}}(\ell)$. We establish this with model-free linear statistics that assume no parametric form for the decay.
@@
Five architectures (ConstGate, SharedGate, DiagGate, GRU, LSTM) $\times$ three optimizers (SGD, AdamW, RMSProp) $\times$ five random seeds =75=75 runs. Hidden size H=64H{=}64, sequence length T=500T{=}500, batch size B=256B{=}256, 200 training epochs on a sinusoidal delayed-regression task. For each run, 800 $(t,\ell)$ samples per lag are drawn at 12 lags spanning 1–245. At every sample we compute the exact envelope $f_{\mathcal{M}}(\ell)$ and the first-order approximation $f(\ell)$ in Eq. 8. We report two statistics: the Spearman rank correlation $\rho$, which tests whether the two envelopes rank lags in the same order, and the Pearson correlation $r$ on the $\log_{10}$-transformed envelopes, which tests whether the decay shapes are linearly related in log-space. Both are computed on only 12 points (one per lag), making them deliberately strict: even a single outlier substantially reduces the correlation.
@@ Results.
Spearman $\rho\geq 0.972$ across all 75 runs confirms that the first-order expansion never disrupts the lag ordering of the envelope. Pearson $r\geq 0.823$ (mean $0.95$) on the $\log_{10}$-transformed envelopes establishes that $\log f\approx a\,\log f_{\mathcal{M}}+b$, i.e. the two envelopes are related by a power-law rescaling $f\propto f_{\mathcal{M}}^{\,a}$. Such a rescaling preserves the monotonic and scaling structure on which the learnability analysis depends.
@@ Appendix D Projected adaptive base rate via Rayleigh quotient
Appendix E Noise floor of the averaged matched statistic
@@
This appendix clarifies the assumptions underlying the noise model used in the manuscript and justifies the scaling of stochastic fluctuations in the averaged matched statistic. The key distinction is between exact finite-sample distributions and domain-of-attraction arguments, which suffice to characterize scaling behavior. Throughout, we rely on standard properties of heavy-tailed random variables [42].
By construction, the averaged matched statistic is the cross-sequence average $\widehat{S}_{N}(\ell)$ defined in Eq. (17), built from the sequence-level averages $\bar{S}^{(n)}_{\ell}$ in Eq. (16). Independence across sequences implies that the sequence-level summands $\bar{S}^{(n)}_{\ell}$ are independent for fixed $\ell$. A raw-signed empirical proxy can be defined with the same two-stage averaging structure by replacing the oracle sign-oriented summand by its raw signed analogue, as discussed in the experiments.
Note that, while the theoretical formulation of the matched statistic in Eq. (13) involves nonlinear operations (the oracle sign factors $\operatorname{sgn}(m_{q}(\ell))$), the practical statistic $\widetilde{S}_{t,\ell}$ in Eq. (31) is a purely linear function of the gradient: a weighted sum over neurons with weights given by the generalized effective learning rates $\mu^{(q)}_{t,\ell}$. Empirical studies [48] of stochastic gradients in deep neural networks consistently report heavy-tailed fluctuations, typically with tail index $\alpha\in(1,2)$. Since the practical matched statistic aggregates gradient information through linear operations, this literature motivates modeling its centered sequence-level version $\bar{S}^{(n)}_{\ell}-\mathbb{E}[\bar{S}^{(n)}_{\ell}]$ as belonging to the domain of attraction of a $\mathcal{S}\alpha\mathcal{S}$ law with lag-dependent scale parameter $\sigma_{\alpha}(\ell)$. The detection framework developed in the manuscript depends only on the tail behavior of the matched statistic itself, not on the precise mechanism by which heavy tails arise. This assumption does not require exact $\mathcal{S}\alpha\mathcal{S}$ behavior at finite sample sizes; it only requires regularly varying tails with index $\alpha$. The resulting noise-floor model for the averaged matched statistic $\widehat{S}_{N}(\ell)$ is
@@
$\widehat{S}_{N}(\ell)-\mathbb{E}[\widehat{S}_{N}(\ell)]\overset{\mathrm{model}}{\sim}\mathcal{S}\alpha\mathcal{S}\!\left(0,\sigma_{\alpha}(\ell)N^{1/\alpha-1}\right)=\mathcal{S}\alpha\mathcal{S}\!\left(0,\frac{\sigma_{\alpha}(\ell)}{N^{\,1-1/\alpha}}\right).$ (71)
@@
This display should be read as the canonical stable surrogate associated with the domain-of-attraction assumption above, not as an exact finite-sample identity. Note that the $N^{-1/\alpha}$ scaling applies to sums of $\alpha$-stable variables, whereas the empirical average introduces an additional factor $1/N$, yielding the rate $N^{-(1-1/\alpha)}$. Thus, the typical magnitude of stochastic fluctuations decays as $N^{-(1-1/\alpha)}$. For $\alpha=2$ this recovers the classical Gaussian rate $N^{-1/2}$, while for $\alpha<2$ concentration is slower due to heavy-tailed noise. This noise floor determines the detectability thresholds and sample-complexity bounds derived in the manuscript and underlies the compression of the learnability window in the heavy-tailed regime.
@@ Appendix F LAN-based KL bound for averaged $\alpha$-stable location models
For completeness, we derive the KL lower bound used in Section 5.4.2 for the averaged matched statistic $\widehat{S}_{N}(\ell)$ defined in Eq. (17). Unlike the classical LAN setting for product measures, the present analysis concerns a shrinking-scale $\alpha$-stable location family. The argument follows the general LAN framework for triangular arrays of statistical experiments with vanishing noise levels [51].
@@
We consider the statistical experiment induced by the averaged matched statistic $\widehat{S}_{N}(\ell)$ at a fixed lag $\ell$, where each sequence contributes the temporally averaged scalar $\bar{S}^{(n)}_{\ell}$ from Eq. (16). As shown in Appendix E, under heavy-tailed fluctuations with tail index $\alpha>1$, the centered statistic $\widehat{S}_{N}(\ell)-\mathbb{E}[\widehat{S}_{N}(\ell)]$ exhibits stochastic fluctuations whose typical magnitude scales as $N^{-(1-1/\alpha)}$.
To formalize hypothesis testing at lag $\ell$, we model the distribution of the centered averaged matched statistic by a location family with shrinking scale. Specifically, let $Z$ be a $\mathcal{S}\alpha\mathcal{S}$ random variable with unit scale and characteristic exponent $\alpha>1$, and define the observation
@@
Here, $Y_{N}$ represents the random fluctuation of the averaged matched statistic, and $\theta\in\mathbb{R}$ is a location parameter encoding the presence or absence of a lag-$\ell$ dependency.
@@
For $\alpha>1$, $\mathcal{S}\alpha\mathcal{S}$ laws are absolutely continuous and admit a density. Denoting by $P_{\theta,N}$ the law of $Y_{N}$, its density is given by
@@
where $f_{0}$ is the density of a fixed $\mathcal{S}\alpha\mathcal{S}$ distribution with unit scale.
@@
For $\alpha>1$, $\mathcal{S}\alpha\mathcal{S}$ densities possess finite Fisher information for the location parameter [23, 33, 42, 51]. As $s_{N}\to 0$, the family $\{P_{\theta,N}\}_{N\geq 1}$ forms a triangular array of location models that is locally asymptotically normal with normalization rate
@@
In the main text, the detection and non-detection hypotheses correspond to the two location shifts in Eq. (18), namely $\theta_{1}=+\tfrac{1}{2}\Delta(\ell)$ and $\theta_{0}=-\tfrac{1}{2}\Delta(\ell)$, so that $\theta_{1}-\theta_{0}=\Delta(\ell)$.
The averaged matched statistic has a shrinking noise scale $s_{N}=\sigma_{\alpha}(\ell)N^{1/\alpha-1}$ (Appendix E). To express the fixed separation in the LAN framework, we measure it in units of this scale and define the normalized separation
@@ Appendix G From KL divergence to mutual information and Fano bounds
Fix a lag $\ell$. The hypotheses in Sec. 5.4.2 correspond to two $\alpha$–stable location models for the finite-sample matched statistic $S=\widehat{S}_{N}(\ell)$ from Eq. (17). Introduce a binary label